OpenStudy (anonymous):

F(x)=x^2+6 and G(x)=x+8\x find (g o f) (-7) help pleaaaaase (:

OpenStudy (anonymous):

first find \[f(-7)\] do you know that one?

OpenStudy (anonymous):

noo, well, i think sooo...i really don't quit understand this stuf :/ do i plug it in?1

OpenStudy (anonymous):

yes, replace x by -7 in f(x)

OpenStudy (anonymous):

\[F(-7)=(-7)^2+6=49+6=55\]

OpenStudy (anonymous):

now to find \[g(F(-7))=g(55)\] replace x by 55 in g(x)

OpenStudy (anonymous):

soo i do 55 * -7?

OpenStudy (anonymous):

it is not clear to me what g(x) is. did you write \[g(x)=\frac{x+8}{x}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

g(x) X+8 over x in fraction form